OpenStudy (anonymous):

Solve This Equation. 4c+2c+6=24

OpenStudy (anonymous):

Just sum sames ;) Sames are 4c and 2c there ;) If you sum them have : (4+2)c+6=24 Got it ?;) By the way ... Welcome in open study ;)!!!

OpenStudy (anonymous):

@E.ali

OpenStudy (anonymous):

i dont get it

OpenStudy (anonymous):

Same sentences : Same sentence are the sentence we can sum them for example :2x and 8x The same should : have : 1 . same words : x and x y and y 2 . same power : x^2 and x^2 8x^2 and x^2 Got it ?

OpenStudy (anonymous):

yes i got that

OpenStudy (anonymous):

Very good and smart ;) So look at this example : If we want sum the sentence for example 6x and 8x and 2y first search are they same or not . So 6x and 8x are but 2y not caz they are x and that s y so cant come and sum with them So we have : 6x+8x+2y we just sum : 6x and 8x and just write 2y : 6x+8x+2y=14x+2y Just this ;) Got it ?:)

OpenStudy (anonymous):

Yes :)

OpenStudy (anonymous):

hey beautifulthang do u want to find c?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

okkk then u do it this way: 4c+2c+6=24 6c=24-6 6c=18 c=18/6 c=3 so the answer is 3

OpenStudy (anonymous):

THANK YOU

OpenStudy (anonymous):

@khadeeja : thank you . I forgot to do that ...:P And sorry @_beautifulthang

OpenStudy (anonymous):

@E.ali its okay !

OpenStudy (anonymous):

Thant was my task ... and thank you ;)